I can't seem to find the solution to this problem. I would appreciate some guidance. Thank you. A solution containing 30% insecticide is to be mixed with a solution containing 50% insecticide to make 200 L of a solution containing 42% insecticide. How much of the 50% insecticide solution should be used? liters
Answer by scott8148(5880)

let x="liters of 50%", so 200-x="liters of 30%"

x(50%)+(200-x)(30%)=200(42%) __ .5x+60-.3x=84 __ .2x=24